1. You are in the market for a new car, and you have narrowed the search to a Tesla with a price of $38,000 and will save you $500 in gasoline costs for each of the next five years and a GMC for $35,600. Assuming the opportunity cost is 3%, which car should you purchase if the goal is to spend the less amount of money for the car and gas over the next five years?
Answer #1

ANSWER:

We need to find the present worth of the tesla car.

i = 3% and n = 5 years

pw of tesla car = purchase price - savings per year(p/a,i,n)

pw of tesla car = 38,000 - 500(p/a,3%,5)

pw of tesla car = 38,000 - 500 * 4.58

pw of tesla car = 38,000 - 2,290

pw of tesla car = 35,710

pw of gmc car = $35,600

sine the pw of gm car is less then that of tesla car , therefore i will purchase the gmc car.